# Video-Language Alignment Pre-training via Spatio–Temporal Graph Transformer

### v1

@@ Abstract
Video-language alignment is a crucial multi-modal task that benefits various downstream applications, e.g., video-text retrieval and video question answering. Existing methods either utilize multi-modal information in video-text pairs or apply global and local alignment techniques to promote alignment precision. However, these methods often fail to fully explore the spatio-temporal relationships among vision tokens within video and across different video-text pairs. In this paper, we propose a novel Spatio–Temporal Graph Transformer module to uniformly learn spatial and temporal contexts for video-language alignment pre-training (dubbed STGT). Specifically, our STGT combines spatio-temporal graph structure information with attention in transformer block, effectively utilizing the spatio-temporal contexts. In this way, we can model the relationships between vision tokens, promoting video-text alignment precision for benefiting downstream tasks. In addition, we propose a self-similarity alignment loss to explore the inherent self-similarity in the video and text. With the initial optimization achieved by contrastive learning, it can further promote the alignment accuracy between video and text. Experimental results on challenging downstream tasks, including video-text retrieval and video question answering, verify the superior performance of our method. Code is available at: https://github.com/GXYM/STGT.
@@
Video-language alignment, Spatio–temporal graph, Self-similarity alignment, Video-text retrieval, Video question answering.
@@ 1 Introduction
Recently, image-language pre-training has achieved significant success in cross-modal representation learning [1, 2, 3]. Adapting a robust image-language pre-trained model for video-language pre-training can offer significant advantages by leveraging knowledge derived from images. Some methods [4, 5, 6] directly applied image-language pre-trained models to video-text tasks, still outperforming video-specific models. As an essential multi-modal task in vision-language pre-trained models, video-language alignment can benefit various tasks, such as video-text retrieval [4, 7] and video question answering [8, 9].
@@
Some methods [5, 6, 10] have adopted image-language pre-training models for video-language pre-training. CLIP4Clip [5] transfers knowledge from pre-trained CLIP [11] for video clip retrieval. It adopts mean-pooling to compute similarity for averaging multiple frame features extracted from the pre-trained CLIP model, and re-trains on the the large-scale video-language Howto100M [12] dataset. However, only averaging or fusing video frame features may not sufficiently capture the temporal dynamics and contextual information in videos. X-CLIP [7] solves this problem by introducing a cross-frame communication transformer (CCT) and a multi-frame integration transformer (MIT). The CCT uses frame-level features from a pre-trained image-language model to capture semantics and model dependencies between frames, while MIT transfers frame-level features to video-level seamlessly. However, these strategies have yet to be fully validated in tasks such as video-text retrieval and video question answering. Although some approaches have tried to capture temporal dynamics and inter-frame dependencies, exploring the spatio-temporal relationships between vision tokens to boost feature representation capability still remain to be improved.
@@
Another problem is that the similarity relationship between different video-text pairs remains under-explored. Traditional contrastive learning based methods [7, 13, 4] mainly focuses on the matching relationship within annotated video-text pairs. In these methods, video-text pair A is deemed a positive sample, while pair B is considered a negative sample for A, as shown in Fig. LABEL:fig:intro1 (c). The contrastive loss aims to maximize the similarity between the embeddings of video A and text A while minimize the similarity between the embeddings of video A and text B, as indicated by the red line in Fig. LABEL:fig:intro1 (c). However, it does not explicitly consider relationships between different videos or texts. In Fig. LABEL:fig:intro1 (c), videos A and B have similar scene and content (similarity score is 0.8). However, texts A and B exhibit a lower similarity of 0.3 due to the differences in observation perspectives and language expression. Despite the low similarity between text A and B, text A can still describe video B and vice versa, due to their high visual similarity. In such cases, it’s unnecessary to force the similarity of learned embeddings between video A and text B, or learned embeddings between video B and text A to approach zero. Therefore, designing an objective loss that appropriately assigns similarity relationships to these cross video-text pairs is a valuable research problem.
@@
To address these problems, we propose a novel spatio–temporal graph transformer network for video-language alignment. Specifically, the spatio–temporal graph transformer module (dubbed STGT) combines the graph and transformer models to effectively and uniformly learn videos’ spatial and temporal features. By measuring the similarity between vision tokens in video frames, we construct a spatio-temporal graph containing both temporal and spatial contextual information, as shown in Fig. LABEL:fig:intro1 (a). Then, the STGT directly integrates the spatio-temporal graph’s topology and edge weights into the attention of the transformer block, effectively leveraging the spatio-temporal context offered by the graph structure. In this way, our method can model the relationships between vision tokens, promoting the precision of video-text alignment for benefiting downstream tasks. Additionally, we propose a self-similarity alignment loss to explore the inherent self-similarity in the video and text. With the initial optimization achieved by contrastive learning, it can further promote the alignment accuracy between video and text. Experimental results on challenging downstream tasks, including video-text retrieval and video question answering, verify the superior performance of our method. In summary, the main contributions of this paper are three-fold:
@@
  • We propose a novel spatio–temporal graph transformer module to explore spatial and temporal relationships for video-language alignment uniformly.
@@
  • We propose a self-similarity alignment loss to explore the inherent self-similarity in the video and text, further promoting the accuracy of video-text alignment.
@@
  • Experiments on challenging downstream tasks, including text-video retrieval and video question answering, verify the superior performance of our method.
@@ 2.1 Vision-Language Pre-training
Vision-language pre-training aims to develop a unified multi-modal representation, enhancing performance on various downstream tasks, such as video-text retrieval and video question answering. The dual-encoder is a common approach in visual-language alignment. Methods [14, 11, 5, 15, 4, 16, 7, 17, 18] employ two separate encoders to independently extract features for visual and textual data. CLIP [11] effectively applies contrastive learning to learn image-language alignment from a large volume of noisy image-text pairs, achieving remarkable performance on vision-language tasks, as demonstrated in [19, 20, 21]. Vita-CLIP [6] introduces a multi-modal prompt learning scheme that balances supervised and zero-shot performance under a single unified training framework. Cross-fusion based methods [22, 23, 24, 25, 26, 27] use a cross-modal fusion encoder to enhance the interactions between vision and text features. BEiT-3 [28] and VLMo [29] employ modality-specific feed-forward networks and a shared self-attention layer in each block to facilitate flexible single-modality learning and cross-modal interactions. In ALPRO [30], the authors introduce a video-text contrast (VTC) loss to align instance-level unimodal video-text features and design a prompt entity module to learn fine-grained alignment between visual regions and textual entities in a self-supervised manner. Following [1], BLIP-2 [31] integrates pre-trained visual models with frozen parameters and large-scale language models.
@@
Figure 1: Overview of the proposed framework. Firstly, vision tokens are obtained by a pre-trained ViT. Subsequently, a spatio-temporal graph transformer module is used to enhance global and local features. Finally, a self-similarity alignment loss is implemented to optimize video-text alignment following contrastive learning optimization.
2.2 Video-Text Retrieval
@@
Previous research [32, 33] in video-text retrieval has primarily focused on developing intricate fusion mechanisms to facilitate cross-modal learning. Recently, vision-language pre-training models [4, 7, 5, 34, 35, 36, 37, 38, 39] have made significant strides in zero-shot and fine-tuned video-text retrieval tasks. Clip4Clip [5] introduces a similarity calculator to directly transfer the robust knowledge from the pre-trained CLIP and continue pre-training on a large-scale video-language dataset for video clip retrieval. VideoClip [4] trains a transformer for video and text by contrasting temporally overlapping positive video-text pairs with hard negatives from nearest neighbor retrieval. In DRL [10], considering the inherent sequential structure in both text and video inputs, a weighted token-wise interaction module is implemented to decouple the content and adaptively exploit the pair-wise correlations. In [40], the authors introduce a text-video learning framework with progressive spatio-temporal prototype matching, which can uncover semantic diversity in videos for dynamic matching.
2.3 Video Question Answering
@@
Video question answering involves automatically generating responses to questions based on video context. This task necessitates a thorough understanding of both the video content and the language used in the questions. To extract more reliable multi-modal representations for the video question answering task, some methods [41, 42, 43, 44, 45] concentrate on enhancing spatio-temporal attention networks, while others [8, 9, 46] focus on designing superior question-video relation networks. Recently, the multi-modal encoder with video-language pre-training [47, 13] has been employed to perform token-level cross-modal fusion. Clover [13] uses a tri-modal alignment pre-training task to improve video-language feature alignment, and the pre-trained model is then applied to the video question answering task. VideoChat [48] designs an end-to-end chat-centric video understanding system by pre-training a vision transformer (ViT) and a large language model (LLM).
@@ 3.1 Overview
We use a pre-trained vision transformer (ViT) as the vision encoder and a pre-trained BERT from BLIP-2 [31] as the text encoder. As shown in Fig. 1, our method consists of three main components. Firstly, we extract vision tokens using ViT and add temporal and 2D spatial embeddings to enhance spatio-temporal relationships. Then, a spatio-temporal graph transformer module refines global and local features. This module constructs a spatio-temporal graph based on local token similarity and integrates the graph’s topology and edge weights into attention for leveraging spatio-temporal context effectively. Additionally, a cross-frame attention block enhances cross-frame communication of global information. Finally, we propose a self-similarity alignment loss to optimize video-language alignment, supplementing contrastive loss for improving accuracy and performance.
@@ 3.2 Vision Feature Encoder
Besides spatial encoding, we apply a set of learnable temporal embedding $\{T_{1},...,T_{i},...,T_{m}\}$ to vision tokens. This embedding contain temporal information, which allows the model to encode contextual information over time, thereby understanding the temporal dynamics in video. Both embeddings (2D spatial embedding $E_{2p}$ and temporal embedding $E_{t}$) are then combined with vision tokens by element-wise addition, denoted as:
@@
where $Tk_{i}(j)$ represents the i-$th$ vision token of the j-$th$ frame in video; $F_{v}(j)$ represents the features of the j-$th$ frame in video. The range of $i$ is from $1$ to $n^{2}$, and the range of $j$ is from $1$ to $m$. The combined embedding $F_{v}(j)$ incorporates the vision tokens, 2D spatial embedding $E_{2p}$, and temporal embedding $E_{t}$. This strategy allows the spatio-temporal graph transformer module to access both raw visual data and spatial-temporal context. By linking different tokens, the model learns a context-rich representation of each tokens, thereby enhancing its understanding of the video sequence.
@@ 3.3 Spatio–Temporal Graph Transformer
Global Features. The $\mathbf{[CLS]}$ vision token is important as it holds the global features ($\mathcal{X}_{g}$) of each video frame. These features give an important global representation of video frame. In Section 3.2, we use temporal embeddings to encode the time position of frame in video sequence. By merging these with global features, our method considers the video’s temporal attributes. To enable the exchange of global information across frames, we employ a cross-frame attention block, as
@@
In videos, the observed similarity principle extends beyond individual frames to include multiple frames. By effectively utilizing these similarities, local features can be thoroughly enhanced. In our method, we construct a spatio-temporal graph ($g(\mathcal{X}_{l},\mathcal{A})$) based on the similarities of local tokens. Each local vision token is regarded as a node in the graph. Consequently, the local feature matrix can directly represent the node feature matrix ($\mathcal{X}_{l}=flatten([F_{v}(1),...,F_{v}(i),...,F_{v}(m)],(0,1))$) of the constructed graph. These features integrate both the positional and temporal information of the tokens, enabling comprehensive spatio-temporal reasoning within the graph. The topology of the spatio-temporal graph can be represented by adjacency matrix $\mathcal{A}$. The adjacency matrix $\mathcal{A}$ defines the connections between nodes in the graph, and contain the spatio-temporal relationships between tokens across different video frames.
@@
Since the similarities between local tokens within frames and between adjacent frames are easily obtained, we can use this information to determine the connections in the spatio-temporal graph. To get the adjacency matrix $\mathcal{A}$, we can iterate over the all nodes in graph and compute their similarities, as
@@
where ||||| | indicates a concatenation operation. By residual transformer block, we effectively integrate both global and local information to capture the synergistic relationship between global and local contexts, thereby enhancing tge comprehensive understanding of the video. To reduce computational complexity, we employ max-pooling to sample the local features, as
@@
3.4 Self-Similarity Alignment Loss
In traditional contrastive learning, only the video-text pairs with the same annotation are considered as positive samples, while other video-text pairs are treated as negative samples. As shown in Fig. 4, the video-text pair A is regarded as a positive sample, while pair B is regarded as a negative sample for A. The contrastive loss aims to maximize the similarity between the embeddings of video A and text A (toward one), while minimize the similarity between the embeddings of video A and text B, as shown in Fig. 4 with red line.
Traditional contrastive learning focuses on the alignment of video and text within the same annotated pairs, but overlooks relationships across different pairs. In Fig. 4, videos A and B have similar scene and content ( similarity score is 0.8). However, texts A and B exhibit a lower similarity of 0.3 due to the differences in observation perspectives, language expression and content summarization. Despite the low similarity between text A and B, text A can still describe video B and vice versa, due to their high visual similarity. In such cases, it’s unnecessary to force the similarity of learned embeddings between video A and text B, or learned embeddings between video B and text A to approach zero. For cross video-text pairs, new similarity relationships need to be considered, as indicated by blue line in Fig. 4.

### v2

# Video-Language Alignment via Spatio–Temporal Graph Transformer

@@ Abstract
Video-language alignment is a crucial multi-modal task for various downstream applications,e.g., video-text retrieval and video question answering. Existing methods either utilize multi-modal information in video-text pairs or apply global and local alignment techniques to promote alignment precision. However, these methods often fail to fully explore the spatio-temporal relationships among vision tokens within video and across different video-text pairs. In this paper, we propose a novel Spatio–Temporal Graph Transformer module to uniformly learn spatial and temporal contexts for video-language alignment (dubbed STGT). Specifically, our STGT combines spatio-temporal graph structure information with attention in the transformer block to explore the spatio-temporal contexts fully. This way, we can effectively model the relationships between vision tokens to promote video-text alignment precision, greatly benefiting downstream tasks. In addition, we propose a novel cross-similarity alignment loss (CSAL) by evaluating the corresponding two video-video and text-text pairs to explore the inherent similarity after the initial optimization achieved by contrastive learning, which can further promote the aligning accuracy. Experimental results on challenging downstream tasks, including video-text retrieval and video question answering, verify the superior performance of our method. The code is available at: https://github.com/GXYM/STGT.
@@
Video-language alignment, Spatio–temporal graph, Cross-similarity alignment, Video-text retrieval, Video question answering.
@@ 1 Introduction
Recently, image-language pre-training has achieved significant success in cross-modal representation learning [1, 2, 3]. Adapting a robust image-language pre-trained model for video-language pre-training can offer significant advantages by leveraging knowledge derived from images. Some methods [4, 5, 6] directly applied image-language pre-trained models to video-text tasks, still outperforming video-specific models. As the critical task in vision-language pre-trained models, the power of video-language alignment can greatly determine the performance of various downstream tasks, such as video-text retrieval [4, 7] and video question answering [8, 9].
@@
Some methods [5, 6] rely on pre-trained image-language models for video-language alignment training. CLIP4Clip [5] transfers knowledge from pre-trained CLIP [10] for video clip retrieval. It adopts mean-pooling to compute similarity for averaging multiple frame features extracted from the pre-trained CLIP model and re-trains on the large-scale video-language Howto100M [11] dataset. However, only averaging or fusing video frame features may not sufficiently capture the temporal dynamics and contextual information in videos. X-CLIP [7] solves this problem by introducing a cross-frame communication transformer (CCT) and a multi-frame integration transformer (MIT). The CCT uses frame-level features from a pre-trained image-language model to capture semantics and model dependencies between frames, while MIT transfers frame-level features to video-level seamlessly. However, these strategies have yet to be fully validated in tasks, such as video-text retrieval and video question answering. Although some approaches have tried to capture temporal dynamics and inter-frame dependencies, exploring spatio-temporal relationships between vision tokens to boost feature representation capability remains to be improved.
@@
Another problem is that the similarity relationship between different video-text pairs still needs to be further explored. Traditional contrastive learning based methods [7, 12, 4] mainly focus on the matching relationship within annotated video-text pairs. In these methods, video-text pair A is deemed a positive sample, while pair B is considered a negative sample for A, as shown in Fig. LABEL:fig:intro1 (c). The contrastive loss aims to maximize the similarity between the embeddings of video A and text A while minimizing the similarity between the embeddings of video A and text B, as indicated by the red line in Fig. LABEL:fig:intro1 (c). However, it does not explicitly consider relationships between different videos or texts. In Fig. LABEL:fig:intro1 (c), videos A and B have similar scenes and content (similarity score is 0.8). However, texts A and B exhibit a lower similarity of 0.3 due to the differences in observation perspectives and language expression. Despite the low similarity between text A and B, text A can still describe video B and vice versa due to their high visual similarity. In such cases, it’s unnecessary to force the similarity of learned embeddings between video A and text B, or learned embeddings between video B and text A to approach zero. Therefore, designing an objective loss that appropriately assigns similarity relationships to these cross video-text pairs is valuable for video-language alignment.
@@
To address these problems, we propose a novel spatio–temporal graph transformer network for video-language alignment. Specifically, the spatio–temporal graph transformer module combines the graph and transformer models to effectively and uniformly learn videos’ spatial and temporal features. By measuring the similarity between vision tokens in video frames, we construct a spatio-temporal graph containing both temporal and spatial contextual information, as shown in Fig. LABEL:fig:intro1 (a). Then, the STGT directly integrates the spatio-temporal graph’s topology and edge weights into the attention of the transformer block, effectively leveraging the spatio-temporal context offered by the graph structure. In this way, our method can model the relationships between vision tokens, promoting the precision of video-text alignment for benefiting downstream tasks. Additionally, we propose a novel cross-similarity alignment loss by evaluating the corresponding two video-video and text-text pairs to explore the inherent similarity after the initial optimization achieved by contrastive learning, which can further promote the aligning accuracy. Experimental results on challenging downstream tasks, including video-text retrieval and video question answering, verify the superior performance of our method. In summary, the main contributions of this paper are three-fold:
@@
  • We propose a novel spatio–temporal graph transformer module (STGT) that combines spatio-temporal graph structure information to fully explore spatial-temporal contexts between video and text pairs for video-language alignment.
@@
  • We propose a novel cross-similarity alignment loss (CSAL) to explore the inherent self-similarity via evaluating the corresponding two video-video and text-text pairs, further promoting the accuracy of video-text alignment.
@@
  • Experiments on challenging downstream tasks, e.g. text-video retrieval and video question answering, verify the superior performance of our method.
@@ 2.1 Vision-Language Pre-training
Vision-language pre-training aims to develop a unified multi-modal representation, enhancing performance on various downstream tasks, such as video-text retrieval and video question answering. The dual-encoder is a common approach in visual-language alignment. Methods [13, 10, 5, 14, 4, 15, 7, 16, 17, 18] employ two separate encoders to independently extract features for visual and textual data. CLIP [10] effectively applies contrastive learning to learn image-language alignment from a large volume of noisy image-text pairs, achieving remarkable performance on vision-language tasks, as demonstrated in [19, 20, 21, 22, 23]. In VATT [15], the authors employ contrastive learning to align the videos, audios and texts, and achieve impressive performance on the downstream tasks. Vita-CLIP [6] introduces a multi-modal prompt learning scheme that balances supervised and zero-shot performance under a single unified training framework.
Cross-fusion based methods [24, 25, 26, 27, 28, 29] use a cross-modal fusion encoder to enhance the interactions between vision and text features. BEiT-3 [30] and VLMo [31] employ modality-specific feed-forward networks and a shared self-attention layer in each block to facilitate flexible single-modality learning and cross-modal interactions. ALBEF [26] incorporates the image and text features into a cross-modal attention-based encoder to generate fused features. In ALPRO [32], the authors introduce a video-text contrast (VTC) loss to align instance-level unimodal video-text features and design a prompt entity module to learn fine-grained alignment between visual regions and textual entities in a self-supervised manner. Following [1], BLIP-2 [33] integrates pre-trained visual models with frozen parameters and large-scale language models.
@@
Figure 1: Overview of the proposed framework. Firstly, vision tokens are obtained by a pre-trained ViT. Subsequently, a spatio-temporal graph transformer module is used to enhance global and local features. Finally, a self-similarity alignment loss is implemented to optimize video-text alignment following contrastive learning optimization. The BLIP-2 [33] attention masking strategy for each objective to control query-text interaction.
2.2 Downstream Tasks
@@
Video-Text Retrieval. Previous research [34, 35] in video-text retrieval has primarily focused on developing intricate fusion mechanisms to facilitate cross-modal learning. Recently, vision-language pre-training models [4, 7, 5, 36, 37, 38, 39, 40, 41] have made significant strides in zero-shot and fine-tuned video-text retrieval tasks. Clip4Clip [5] introduces a similarity calculator to directly transfer the robust knowledge from the pre-trained CLIP and continue pre-training on a large-scale video-language dataset for video clip retrieval. VideoClip [4] trains a transformer for video and text by contrasting temporally overlapping positive video-text pairs with hard negatives from nearest neighbor retrieval. In DRL [42], considering the inherent sequential structure in both text and video inputs, a weighted token-wise interaction module is implemented to decouple the content and exploit the pair-wise correlations. In [43], the authors introduce a text-video learning framework with progressive spatio-temporal prototype matching, which uncover semantic diversity in videos for dynamic matching.
@@
Video Question Answering. Video question answering involves automatically generating responses to questions based on video context. This task necessitates a thorough understanding of both the video content and the language used in the questions. To extract more reliable multi-modal representations for the video question answering task, some methods [44, 45, 46, 47, 48] concentrate on enhancing spatio-temporal attention networks, while others [8, 9, 49] focus on designing superior question-video relation networks. Recently, the multi-modal encoder with video-language pre-training [50, 12] has been employed to perform token-level cross-modal fusion. Clover [12] uses a tri-modal alignment pre-training task to improve video-language feature alignment, and the pre-trained model is then applied to the video question answering task. VideoChat [51] designs an end-to-end chat-centric video understanding system by pre-training a vision transformer (ViT) and a large language model (LLM).
@@ 3.1 Overview
We use a pre-trained vision transformer (ViT) as the vision encoder and a pre-trained BERT from BLIP-2 [33] as the text encoder. As shown in Fig. 1, our method consists of three main components. Firstly, we extract vision tokens using ViT and add temporal and 2D spatial embeddings to enhance spatio-temporal relationships. Then, a spatio-temporal graph transformer module refines global and local features. This module constructs a spatio-temporal graph based on local token similarity and integrates the graph’s topology and edge weights into attention for leveraging spatio-temporal context effectively. Additionally, a cross-frame attention block enhances cross-frame relationship of global information. Finally, we propose a cross-similarity alignment loss to optimize video-language alignment, supplementing contrastive loss for improving performance.
@@ 3.2 Vision Feature Encoder
Besides spatial encoding, we apply a set of learnable temporal embedding $\{T_{1},...,T_{i},...,T_{m}\}$ to vision tokens. This embedding contains temporal information, which allows the model to encode contextual information over time, thereby understanding the temporal dynamics in video. Both embeddings (2D spatial embedding $E_{2p}$ and temporal embedding $E_{t}$) are then combined with vision tokens by element-wise addition, denoted as:
@@
where $Tk_{i}(j)$ represents the i-$th$ vision token of the j-$th$ frame in video; $F_{v}(j)$ represents the features of the j-$th$ frame in video. The range of $i$ is from $1$ to $n^{2}$, and the range of $j$ is from $1$ to $m$. The combined embedding $F_{v}(j)$ incorporates the vision tokens, 2D spatial embedding $E_{2p}$, and temporal embedding $E_{t}$. This strategy allows the spatio-temporal graph transformer module to access both raw visual data and spatial-temporal context. By linking different tokens, the model learns a context-rich representation of each token, thereby enhancing its understanding of the video sequence.
@@ 3.3 Spatio–Temporal Graph Transformer
Global Features. The $\mathbf{[CLS]}$ vision token is important as it holds the global features ($\mathcal{X}_{g}$) of each video frame. These features give an important global representation of the video frame. In Section 3.2, we use temporal embeddings to encode the time position of the frame in a video sequence. By merging these with global features, our method considers the video’s temporal attributes. To enable the exchange of global information across frames, we employ a cross-frame attention block, as
@@
In videos, the observed similarity principle extends beyond individual frames to include multiple frames. By effectively utilizing these similarities, local features can be thoroughly enhanced. In our method, we construct a spatio-temporal graph ($g(\mathcal{X}_{l},\mathcal{A})$) based on the similarities of local tokens. Each local vision token is regarded as a node in the graph. Consequently, the local feature matrix can directly represent the node feature matrix ($\mathcal{X}_{l}=flatten([F_{v}(1),...,F_{v}(i),...,F_{v}(m)],(0,1))$) of the constructed graph. These features integrate both the positional and temporal information of the tokens, enabling comprehensive spatio-temporal reasoning within the graph. The topology of the spatio-temporal graph can be represented by adjacency matrix $\mathcal{A}$. The adjacency matrix $\mathcal{A}$ defines the connections between nodes in the graph, and contains the spatio-temporal relationships between tokens across different video frames.
@@
Since the similarities between local tokens within frames and between adjacent frames are easily obtained, we can use this information to determine the connections in the spatio-temporal graph. To get the adjacency matrix $\mathcal{A}$, we can iterate over all the nodes in the graph and compute their similarities, as
@@
where ||||| | indicates a concatenation operation. By residual transformer block, we effectively integrate both global and local information to capture the synergistic relationship between global and local contexts, thereby enhancing the comprehensive understanding of the video. To reduce computational complexity, we employ max-pooling to sample the local features, as
@@
3.4 Cross-similarity Alignment Loss
In traditional contrastive learning, only the video-text pairs with the same annotation are considered as positive samples, while other video-text pairs are treated as negative samples. As shown in Fig. 4, the video-text pair A is regarded as a positive sample, while pair B is regarded as a negative sample for A. The contrastive loss aims to maximize the similarity between the embeddings of video A and text A (toward one), while minimizing the similarity between the embeddings of video A and text B, as shown in Fig. 4 with a red line.
Traditional contrastive learning focuses on the alignment of video and text within the same annotated pairs, but overlooks relationships across different pairs. In Fig. 4, videos A and B have similar scene and content ( similarity score is 0.8). However, texts A and B exhibit a lower similarity of 0.3 due to the differences in observation perspectives, language expression and content summarization. Despite the low similarity between text A and B, text A can still describe video B and vice versa, due to their high visual similarity. In such cases, it’s unnecessary to force the similarity of learned embeddings between video A and text B, or learned embeddings between video B and text A to approach zero. For cross video-text pairs, new similarity relationships need to be considered, as indicated by the blue line in Fig. 4.